A cylinder has a length of 6.0cm and a curved surface area of 130cm^2. Find its radius.
If you could briefly explain how to do it, give hints or show working, it would be greatly appreciated.
Thank you very much (:

Answer by fcabanski(1391)

The length of the cylinder is usually called its height.


The lateral surface area - the area of the curved surface of a cylinder - is the circumference of the circle formed by the curved surface x the height of the cylinder (in this case what is called the length = 6cm). The circumference is 2*pi*r.


2*pi*r*6 = 130


r = 130/2*pi*6 = approximately 3.45 centimeters.